### v1

# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

Abstract
In the quest to facilitate the deep intelligence of Large Language Models (LLMs) accessible in final-end user-bot interactions, the art of prompt crafting emerges as a critical yet complex task for the average user. Contrast to previous model-oriented yet instruction-agnostic Automatic Prompt Optimization methodologies, yielding polished results for predefined target models while suffering rapid degradation with out-of-box models, we present Free-form Instruction-oriented Prompt Optimization (FIPO). This approach is supported by our large-scale prompt preference dataset and employs a modular fine-tuning schema. The FIPO schema reimagines the optimization process into manageable modules, anchored by a meta prompt that dynamically adapts content. This allows for the flexible integration of the raw task instruction, the optional instruction response, and the optional ground truth to produce finely optimized task prompts. The FIPO preference dataset is meticulously constructed using the optimal and suboptimal LLMs, undergoing rigorous cross-verification by human experts and analytical models. Applying the insights from the data with Tulu2 models and fine-tuning strategies, we validate the efficacy of FIPO schema across five public benchmarks [^1] Codes, data and scripts are here: https://github.com/LuJunru/FIPO_Project..
FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated their exceptional intelligence Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) by excelling in various standard benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). Nevertheless, LLMs can not fully showcase their potential without carefully crafted prompts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b), especially when engaging with non-experts in final-end user-bot interactions.
Figure 1: An example of comparison between other LLM-based discrete APO strategies with our FIPO.
In recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, fpcusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce gradient-free, straightforward strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)). These approaches typically involve iterative processes between a prompt generator LLM and an evaluator LLM. Inspired by GAN Goodfellow et al. (2020), the generator iteratively explores and tests superior prompts, while the evaluator provides feedback on the optimized prompts and the testing results. However, despite their effectiveness, several drawbacks exist: (1) There are unavoidable instability in API-reliant frameworks; (2) The ad-hoc optimization tends to be model-centric, leading to performance degradation when tested with out-of-box models; (3) Many LLM-based discrete APO methods focus on the optimization of instruction-agnostic task descriptions, whereas optimizing instruction-oriented sections is more critical for assisting smaller LLMs.
To address the challenges and break free from the constraints imposed by pre-defined optimization templates in some existing methods Zhou et al. (2023), we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). Figure 1 illustrates a comparison between two mainstream LLM-based discrete APO strategies and our FIPO. On the left-hand is an example of In-Context-Learning Dong et al. (2023) prompt, consisting of a leading introduction sentence, two separate examples, and the testing task instruction “Calculate the average value of the list”. In particular, the two in-context examples follow a consistent format: a one-line question, a line-by-line Chain-of-Thought Wei et al. (2022) solution, and a final answer reporting template (“The answer is X”). It is worth mentioning that such few-shot template, is widely-accepted as the official testing format in popular benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Hendrycks et al. (2021).
In the right-hand section of Figure 1, we present a list of optimized prompts generated through various methods, arranged from top to bottom. The top-right method focuses on optimizing the top leading sentence. We define this category as “Instruction-agnostic” Optimization. In contrast, the middle-right approach, labeled as “Template-based” Optimization, involves incorporating specific guiding sentences into the entire prompt, accopmanied with a series of blank tokens that are filled in to optimize the testing instruction. Our proposed “FIPO” method, listed in the right-bottom part, avoids using preliminary templates and only considers the testing instruction for “Instruction-oriented” Optimization. Among all the optimized prompts, FIPO stands out as the only method capable of enhancing the testing instruction with embedded chain-of-thought guidance, instructing downstream models to “find the sum” and “divide it by the total number of elements” to more effectively complete the “Calculate the average value” task.
Our contributions can be summarized as follows:
  • We introduce our research of free-form instruction-oriented prompt optimization (FIPO), and subsequently devise a modular schema for in-depth investigation.
  • Based on the schema, we compile a large-scale prompt preference dataset, and explore with several mainstream fine-tuning strategies: Supervised Fine-tuning, Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). We further develop an Iterative Preference Learning (IPL) pipeline for self-rewarding preference optimization.
  • We evaluate the efficacy and adaptability of fine-tuned FIPO optimizer across various downstream benchmarks, and obtain significant improvements compared with previous LLM-powered discrete APO methods.
2 Related Work
Two related topics are involved: Automatic Prompt Optimization and Preference Optimization.
Figure 2: Harnessing data from a suboptimal LLM GPT-3.5 and an optimal LLM GPT-4.
2.1 Automatic Prompt Optimization
Automatic Prompt Optimization (APO) is an simple yet effective technique for grabbing potentials of LLMs in various downstream scenarios. Most APO methodologies can be categorized into two types: discrete APO and continuous APO Liu et al. (2023). Discrete APO searches optimized prompts with optimal combinations of discrete tokens Wallace et al. (2019); Shin et al. (2020); Ben-David et al. (2022); Davison et al. (2019); Deng et al. (2022); Zhang et al. (2023); Xu et al. (2022). For instance,  van de Kar et al. (2022) employed text mining for searching candidate prompts from knowledge triplets. While  Yuan et al. (2021),  Haviv et al. (2021) and  Gao et al. (2021) utilized Bart Lewis et al. (2019), Bert Devlin et al. (2018) and T5 Raffel et al. (2019) for optimizing prompts in an paraphrasing manner, respectively.
In contrast, continuous APO proposes to search better prompts in continuous embedding space rather than limited to human-understandable discrete tokens Tsimpoukelli et al. (2021); Zhong et al. (2021); Qin and Eisner (2021); Hambardzumyan et al. (2021); Wang et al. (2023b). Prefix Tuning Li and Liang (2021) and Prompt Tuning Lester et al. (2021) are two well-known continuous APO methods that insert prefix vectors in front of the task sequence, and then update their corresponding parameters. There are also hybrid works to insert continuous embedding into discrete templates Liu et al. (2021); Han et al. (2021).
2.2 Preference Optimization for LLMs
Preference Optimization has been regarded as the key to the final mile of Artificial General Intelligence (AGI) Pan et al. (2023b). Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). A series of following work of DPO are further presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024).
3 Methodology
In this section, we start with task formulation, then introduce our schema, data and training strategies.
3.1 Task Formulation
Named as Free-form Instruction-oriented Prompt Optimization (FIPO), our approach follows the end-to-end text generation process. A neural optimizer model $\bm{M_{o}}$ is tasked with generating an optimized prompt $\bm{\hat{x}^{o}}=\operatorname*{argmax}_{\bm{M_{o}}}p(\bm{\hat{x}^{o}}|\bm{x}^{r},\bm{\hat{y}}^{r},\bm{y}^{r})$, based on the given raw task instruction $\bm{x}^{r}$, raw instruction response $\bm{\hat{y}}^{r}$, and ground truth $\bm{y}^{r}$. The raw instruction response $\bm{\hat{y}}^{r}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{r}|\bm{x}^{r})$ is generated from the raw task instruction $\bm{x}^{r}$ using a neural generator model $\bm{M_{g}}$. The ultimate target is to obtain a superior instruction response $\bm{\hat{y}}^{o}\succeq\bm{\hat{y}}^{r}$, where $\bm{\hat{y}}^{o}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{o}|\bm{\hat{x}}^{o})$. The raw instruction response $\bm{\hat{y}}^{r}$ and ground truth $\bm{y}^{r}$ are only provided during the training phase.
Figure 3: The overall visualization of our fine-tuning strategies.
3.2 Modular Schema and Preference Dataset
In this context, our objective is to develop a modular schema that ensures flexibility in content management, addressing potential exposure bias between the training and the inference phases. Simultaneously, we decide to distill the refined prompt optimization capabilities from prominent yet proprietary LLMs, instead of directly integrating them in an ad-hoc way. Figure 2 illustrates our research schema and how we simply harness contrastive optimized prompts with GPT-3.5 and GPT-4 APIs. We first establish a modular schema as shown in the middle. The meta optimizer prompt takes the raw task instruction $\bm{x}^{r}$ as mandatory input, while considering raw instruction response $\bm{\hat{y}}^{r}$ and ground truth response $\bm{y}^{r}$ as optional. Additional descriptive meta information is then appended for clarity.
The raw task instruction $\bm{x}^{r}$ at the top is sampled from the Alpaca dataset Taori et al. (2023), which contains 52K diverse instructions and corresponding responses $\bm{\hat{y}}^{r}$ generated by the Text-davinci-003 API. We also collect another GPT-4-generated response for the Alpaca dataset from publicly available literature Ghosal et al. (2023). Although there are no official ground truth responses (e.g., from human experts) for the Alpaca data, we consider GPT-4 responses as the ground truth response $\bm{y}^{r}$, given its demonstrated analytical capabilities comparable to humans Pan et al. (2023a). Afterwards, we populate our meta-optimizer prompt with the aforementioned collected instruction $\bm{x}^{r}$, as well as two responses $\bm{\hat{y}}^{r}$ and $\bm{y}^{r}$, and then respectively apply GPT-3.5 and GPT-4 APIs to generate contrastive optimized prompt labels $\bm{x}^{o-}$ and $\bm{x}^{o+}$ in a zero-shot manner. As shown at the bottom of Figure 2, GPT-4 offers a more pedagogical step-by-step optimization compared to GPT-3.5.
We finally narrow down our dataset to 30k samples and report the quality post-checking in Table 1. We adopt cross-validation using three different methods: critiques from an external alignment model UltraRM Cui et al. (2023), self-judgement from the GPT-4, and manual checking by human experts. The “Response” and “Prompt” columns refer to the proportions that GPT-4-generated response and GPT-4-optimized prompt is better than the others, respectively. The average win rates for both categories exceed 85%, ensuring the quality.
GPT-4 Win Rate (%)
Response Prompt Scale
UltraRM 13B Cui et al. (2023) 91.49 82.13 30k
GPT4 Self-check 80.56 92.29 3k
Human Expert 88.29 95.21 1k
Average 86.78 89.88 N/A
Table 1: Quality cross-validation on our dataset.
3.3 Fine-tuning Strategies
We introduce our fine-tuning strategies in Figure 3, consisting of an initial step of dataset diversification followed by strategic fine-tuning.
Dataset Diversification. As illustrated in the left hand of Figure 3, in accordance to our modular schema, we evenly split the original 30k samples as eight types depending on the existence of raw instruction response $\bm{\hat{y}}^{r}$ and ground truth response $\bm{y}^{r}$, as well as a format condition “generation” or “multi-choice”. Robustly fine-tuning neural optimizer models have to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure divergence between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The original responses $\bm{\hat{y}}^{r}$ and $\bm{y}^{r}$ are modified as candidates adhere to the raw task instruction $\bm{x}^{r}$. We therefore set “A” and “B” as new task response and ground truth response [^2] More details of data diversification are in Appendix B..
Strategic Fine-tuning. The right side of Figure 3 introduces several end-to-end fine-tuning strategies that we explored in this work. The right top is the most well-known Supervised Fine-tuning (SFT), which only takes the optimal optimized prompt $\bm{x}^{o+}$ as the supervision signal:
$L_{SFT}(M_{o})=-\mathbb{E}_{(x^{r},\hat{y}^{r},y^{r},x^{o+},x^{o-})\sim D}[\hat{x}^{o}-x^{o+}]^{2}$ (1)
where $D$ stands for the training set.
On the other hand, the right middle shows a very recent yet popular and effective methodology: Preference Optimization, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). Preference Optimization takes contrastive training labels $\bm{x}^{o-}$ and $\bm{x}^{o+}$ as supervision. One of the core differences from Preference Optimization and SFT is the former one not only encourage the generation of optimal preference, but also dampen the generation of suboptimal preference:
$L_{DPO}(M_{o})=-\mathbb{E}_{(x^{r},\hat{y}^{r},y^{r},x^{o+},x^{o-})\sim D}[log\sigma(\beta\Delta)]$ (2)
$L_{IPO}(M_{o})=-\mathbb{E}_{(x^{r},\hat{y}^{r},y^{r},x^{o+},x^{o-})\sim D}[\Delta-\frac{1}{2\beta}]^{2}$ (3)
$\Delta=log\frac{M_{o}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}-log\frac{M_{o}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}$ (4)
where $\bm{\beta}$ is a hyperparameter target. $\bm{M_{ref}}$ refers to the reference model, which is a frozen copy of initial $\bm{M_{o}}$. The equations indicate that IPO is a regularized version of DPO as it limits the optimization range with squares.
Additionally, inspired by recent super-human Reinforcement Learning from AI Feedback (RLAIF) Lee et al. (2023); Anthropic (2022); Yuan et al. (2024), we develop a Iterative Preference Learning (IPL) strategy for self-rewarding preference optimization. After each iteration of preference optimization, we ask the optimizer itself to determine if it successfully generate a superior prompt $\bm{x}^{r+}$ and response $\bm{\hat{y}}^{r+}$, and if so, to automatically replace the previous inferior prompt $\bm{x}^{r}$ and response $\bm{\hat{y}}^{r}$, leading to more rigorous training:
$L_{IPL}(M_{o})=-\mathbb{E}_{(x^{r+},\hat{y}^{r+},y^{r},x^{o+},x^{r})\sim D}G(\Delta)$ (5)
$x^{r+}=M_{o}(x^{r})$ (6)
$x^{r+}=\left\{\begin{array}[]{l}x^{r+},\quad M_{j}(x^{r+},y^{r})\succeq M_{j}(x^{r},y^{r})\\ x^{r},\quad otherwise\end{array}\right.$ (7)
$\hat{y}^{r+}=M_{g}(x^{r+})$ (8)
where $G(*)$ denotes as either IPO or DPO loss style. $\bm{M_{o}}$, $\bm{M_{j}}$ and $\bm{M_{g}}$ refers to the prompt optimizer, the prompt discriminator and the response generator, respectively. However, $\bm{M_{o}}\equiv\bm{M_{j}}\equiv\bm{M_{g}}$ in our settings [^3] Algorithmic introduction of IPL is in Appendix C.. We increment instruction following and discrimination data to maintain its judgement and answering abilities, shown in Appendix B.
The ultimate target of APO lies in the performance enhancement with downstream generators $\bm{M_{g}}$, using final optimized prompt $\bm{\hat{x}^{o}}$ to obtain a superior response $\bm{\hat{y}}^{o}\succeq\bm{\hat{y}}^{r}$. Herein, we evaluate our fine-tuning strategies across a variety of public benchmarks with various downstream generators.
4 Experiments
In this section, we begin with our experimental settings, fine-tuning performances, then follow with comparisons with SOTA, and case analysis [^4] Hyperparameters and training cost are in Appendix A..
4.1 Experimental Settings
Baselines. We compare our results with two SOTA LLM-based discrete APO methods: APE Zhou et al. (2023) and PromptAgent Wang et al. (2023a). APE, stands for Automatic Prompt Engineer, is a template-based strategy that ask one LLM to generate a pool of candidate prompts based on the templates, then select one according to the evaluation scores. PromptAgent eliminates templates and replaces with Monte Carlo Tree Search Abramson (2014) for using a evaluator model to guide the generator. It worth to note that both APE and PromptAgent are training-free, aiming to model-oriented prompt optimization in an ad-hoc manner, with GPT-3.5 and GPT-4 APIs, while we realize completely offline training and inference during the prompt optimization stage. Moreover, since PromptAgent aims to optimized the beginning instruction-agnostic learning sentence, as shown in Figure 1, it requires to attach the following examples and task instruction for few-shot testing. In contrast, APE and our FIPO combine unchanged leading sentence, fixed in-context examples, and optimized task instruction in testing. We use Tulu2 models as our bases, which is a fine-tuned version of Llama2 Touvron et al. (2023) trained on a mix of publicly available datasets Ivison et al. (2023). Three versions of Tulu2 models are released: 7B, 13B and 70B. The “B” here refers to the total number of parameters in unit of billions.
Generation Multi-choice
FIPO Prompt Optimizer GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Human 24.77 36.21 73.35 51.17 51.22 47.79
Best 13B Optimizer 25.38($\uparrow$) 35.93 71.23 49.03 51.18 47.36
SFT-70B (L) 26.44($\uparrow$) 32.46 73.90($\uparrow$) 50.64 50.03 46.33
SFT-70B (F) 21.43 32.92 74.39($\uparrow$) 49.97 51.55($\uparrow$) 46.96
DPO-70B (L) 24.24 41.88($\uparrow$) 76.24($\uparrow$) 55.87($\uparrow$) 54.09($\uparrow$) 51.39($\uparrow$)
DPO-70B (F) 27.74(normal-$\uparrow$) 35.56 74.17($\uparrow$) 54.93($\uparrow$) 52.73($\uparrow$) 49.07($\uparrow$)
IPO-70B (L) 22.95 42.50($\uparrow$) 76.95($\uparrow$) 55.97($\uparrow$) 54.17($\uparrow$) 51.58($\uparrow$)
IPO-70B (F) 25.00($\uparrow$) 39.21($\uparrow$) 76.84($\uparrow$) 56.01($\uparrow$) 54.29($\uparrow$) 50.94($\uparrow$)
IPL-DPO-70B (L) 21.96 42.04($\uparrow$) 75.64($\uparrow$) 56.01($\uparrow$) 55.13($\uparrow$) 51.84($\uparrow$)
IPL-DPO-70B (F) 25.13($\uparrow$) 35.25 74.95($\uparrow$) 50.46 52.12($\uparrow$) 48.10($\uparrow$)
IPL-IPO-70B (L) 21.96 42.92(normal-$\uparrow$) 75.37($\uparrow$) 55.64($\uparrow$) 55.03($\uparrow$) 51.94($\uparrow$)
IPL-IPO-70B (F) 26.67($\uparrow$) 39.60($\uparrow$) 77.11(normal-$\uparrow$) 56.71(normal-$\uparrow$) 56.02(normal-$\uparrow$) 52.13(normal-$\uparrow$)
Table 2: Evaluation results of out-of-box Tulu2-7B model. “Human” refers to use raw few-shot prompts written by humans, while others use optimized prompts generated by different FIPO optimizer. The number attached with the benchmark name refers to the number of in-context examples (e.g., BBH (3) stands for 3-shot evaluation on BBH). The “(L)” or “(F)” letter attached with FIPO optimizer refer to parameter-efficient tuning LORA Hu et al. (2022) and full-parameter tuning, respectively. The “X” part of “IPL-X” refers to the different loss style mentioned in Equation 5. More detailed results are report in Table 8 in Appendix D.
Evaluation Benchmarks. We include five popular benchmarks across two most common formats: (1) GSM8k Cobbe et al. (2021), a generative dataset contains 1.3k primary level math questions; (2) BigBenchHard (BBH) Suzgun et al. (2023), which involves 23 challenging reasoning tasks. BBH has 6.4k testing samples, and asks for generative answering; (3) PiQA Bisk et al. (2020), in which 1.8k common physical knowledge questions are proposed, alongside with binary candidate choices; (4) CosmosQA Huang et al. (2019). There are around 3k commonsense-based multi-choice questions in CosmosQA, equipped with four candidate options; (5) MMLU Hendrycks et al. (2021), which is one of the largest multi-choice benchmarks. MMLU covers 14k questions. For our FIPO experiments, we report results on all five benchmarks. While differently, since both APE and PromptAgent only provide evaluations on 6 tasks of BBH, we report the comparison results aligned with their settings. As for the result metrics, as shown in Figure 1, either “generation” or “multi-choice” benchmarks takes few-shot format with strict answering templates (e.g., “The answer is X”). Herein, we are able to report the accuracy score for all benchmarks.
Generation Multi-choice
Generator Prompt Source GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Llama2-7B Human 8.89 31.21 62.78 43.09 46.58 41.73
Touvron et al. (2023) FIPO Optimizer 11.70 33.50 69.37 52.11 54.56 48.10
Tulu2-13B Human 39.06 36.49 76.62 55.13 57.43 52.53
Ivison et al. (2023) FIPO Optimizer 40.17 40.26 78.58 57.68 59.10 54.79
Baichuan2-13B Human 46.81 37.95 68.56 51.88 57.46 52.36
Yang et al. (2023a) FIPO Optimizer 48.12 39.95 74.77 56.88 58.32 54.35
Table 3: Evaluation results of various downstream generator LLMs.
GPT-3.5 Llama2-7B Tulu2-7B
BBH Tasks Human APE PromptAgent FIPO Human PromptAgent FIPO Human PromptAgent FIPO
Penguins 0.595 0.747 0.797 0.614 0.436 0.426 0.458 0.336 0.470 0.525
Geometry 0.227 0.490 0.670 0.738 0.536 0.544 0.542 0.089 0.350 0.415
Epistemic 0.452 0.708 0.806 0.865 0.611 0.775 0.754 0.581 0.691 0.644
Object Counting 0.612 0.716 0.860 0.660 0.318 0.312 0.350 0.690 0.577 0.637
Temporal 0.720 0.856 0.934 0.756 0.362 0.643 0.683 0.155 0.176 0.210
Causal Judgment 0.470 0.570 0.670 0.759 0.547 0.547 0.616 0.832 0.882 0.891
Average 0.513 0.681 0.790 0.732 0.468 0.541 0.567 0.447 0.524 0.554
Table 4: Comparison between FIPO optimizer and previous methods.
4.2 Experimental Results
Fine-tuning Analysis. We first report main results of testing optimized prompts across downstream benchmarks in Table 2. We fine-tune Tulu2-13B and Tulu2-70B models though different strategies as mentioned above, incorporating with parameter-efficient (e.g., LORA Hu et al. (2022)) and full-parameter way, and then use fine-tuned FIPO optimizer to generate new instruction-oriented prompts for downstream testing. The testing is carried with out-of-box Tulu2-7B model in few-shot format.
We observe several findings: (1) “Small” Tulu2-13B are not up to the difficult prompt optimization task; (2) When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task; (3) When contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%; (4) Optimizer fine-tuned in LORA way achieve better results when combined with DPO, IPO and IPL-DPO strategies. However, full-parameter tuning wins in SFT and IPL-IPO settings; (5) In general, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline. The results indicate the effectiveness of its regularized design in Equation 3; (6) We receive more robust improvements on multi-choice benchmarks. The reason may lie in the output limitations, since evaluations on multi-choice benchmarks are constrained to serial options, while evaluations on generation benchmarks are free to all text.
We further check if our optimized prompts have general gains on different downstream generators in Table 3, and conclude that our optimized instruction-oriented prompts are generally beneficial. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average. The improvements reduced as the generator grows larger. This may be due to the fact that larger LLMs has richer world knowledge, so the enhanced instructions are less necessary.
Analysis on Iterative Preference Learning. To validate the self-rewarding capability of fine-tuning prompt optimizer within our IPL pipeline, we present a further epoch-by-epoch checking in Figure 4. Compared with basic DPO and IPO preference tuning, to pump them into our IPL pipeline can advance to better accuracy expect for the full-parameter tuning of IPL-DPO-70B. It has been visualized that either Lora or full-parameter tunings of DPO under the IPL pipeline brings no self-rewarding gains, showing as marginal fluctuations. However, the IPO fine-tuning fits well. We notice that both Lora and full-parameter tunings of IPO under the IPL pipeline behave raising tendencies, and finally breakthrough the performance ceiling of basic IPO. Particularly, the full-parameter tuning of IPL-IPO-70B exhibits a steeper upward trend.
Figure 4: The tendency visualization of IPL (better view with colors). We denote average accuracy over benchmarks of every checking models as markers. The blue triangles and red triangles refer to IPL-DPO-70B and IPL-IPO-70B models respectively, which is in accordance with the last four lines in Table 2.
Raw Task Instruction:
Calculate the average of the list [12, 34, 56, 75].
Raw Task Response:
Sure, list is a basic data structure in Python. We can compute the average of the list by: (1) list = [12, 34, 56, 75]; (2) average = sum(list) / len(list); (3) print(average). The final answer is sum([12, 34, 56, 75]) / len([12, 34, 56, 75]) = 177 / 4 = 44.25.
Ground Truth Response:
To calculate the average value of a list of numbers, you sum up all the numbers and then divide by the count of numbers in the list. For the list [12, 34, 56, 75], the sum of the numbers is: 12 + 34 + 56 + 75 = 177. There are 4 numbers in the list. Now, divide the sum by the count of numbers to get the average: Average = 177 / 4 = 44.25. So, the average value of the list [12, 34, 56, 75] is 44.25.
Suboptimal Optimized Prompt:
You are an expert in mathematics and someone asks you to calculate the average of a list of numbers. Write a complete solution to explain how to calculate the average step-by-step, using the list [12, 34, 56, 75] as an example. Make sure to provide the final answer in your response.
Optimal Optimized Prompt:
We need to calculate the average of a set of numbers, specifically [12, 34, 56, 75]. Please provide the process in completing the task, including a step-by-step explanation of how to calculate the average, summing all the items in the list, and then dividing this total by the number of elements in the list. Your response should be detailed, clear, and accurate, similar to the manner a math tutor would explain to a student learning this concept for the first time.
Optimized Instruction from SFT-70B (Full)
As an AI language model, your task is to provide a detailed explanation on how to calculate the average of a list of numbers. The list consists of numbers [12, 34, 56, 75]. I would like you to provide a step-by-step guide on how to calculate the average. Please also explain what is meant by “average” in this context, and how the calculation is different from the total sum of the numbers in the list divided by the number of elements in the list. Please start by defining what “average” means in the context of a list of numbers. Explain how this concept relates to the process of calculating the average of a list. Then, provide a step-by-step guide on how to calculate the average of the list [12, 34, 56, 75].
Optimized Instruction from DPO-70B (Full)
Hi AI, could you accurately compute the mean value of this given list of distinct integer numbers? The list consists of: 12, 34, 56, and 75. It would be helpful if you could provide the mathematical formula or steps you took in achieving the solution. Also, make sure to include the final calculated average value in your response.
Optimized Instruction from IPO-70B (Full)
Given the list of integers [12, 34, 56, 75], calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list. Express the result as a floating-point number up to two decimal places. Show the step-by-step calculation to obtain the average and include a brief explanation of the process.
Optimized Instruction from IPL-IPO-70B (Full)
To find the average of the given list of numbers, first, you need to add all the numbers in the list, which are: 12, 34, 56, and 75. Add them up, and you will get a sum of 177. Then, divide the sum by the total number of items in the list, which is 4. So, 177 divided by 4 equals 44.25. Therefore, the average of the list [12, 34, 56, 75] is 44.25.
Table 5: An example from GSM8K benchmark, which is same as the example in Figure 2.
Comparison with SOTA. Except for the effectiveness of fine-tuning strategies, we would like to compare the FIPO research with previous methods. Table 4 reports experimental results on 6 BBH tasks, cross-evaluating with different methods and downstream generators. On the one hand, PromptAgent obtain significant improvements over all tasks, compared with simple few-shot prompts written by humans and optimized prompts generated by APE. However, on the other hand, since PromptAgent runs iterative optimization upon evaluator GPT-4 and generator GPT-3.5, resulting in model-oriented prompt optimization. We thus observe obvious degradation on out-of-box Llama2-7B and Tulu2-7B models. In contrast, our instruction-oriented optimization remains efficacy, and leading the average performances on two small generators.
Case Analysis. In Table 5, we present an example from the downstream GSM8K benchmark for case analysis, which is same as the example shown in Figure 2. We first notice that either Raw Task Response or Ground Truth Response provide correct answers. Nevertheless, the Raw Task Response improperly denotes “list [12,34,56,75]” as a Python data structure. Similarly, both Optimized Prompt generated from GPT APIs provide high-quality paraphrases of the Raw Task Instruction, yet the positive one is superior with chain-of-thought plans (shown with blue in the upper part). We then compare prompt optimization results from our fine-tuned models. The SFT strategy fails with incoherent, misleading clarifications (shown with red in the middle). In contrast, all preference fine-tuned models provide more coherent and concise generations. Particularly, two IPO-based optimizer manages to provide chain-of-thought plans (shown with blue in the bottom), and IPL-IPO-70B further provides cheating notes (noted with underlines).
5 Conclusion
We present the Free-form Instruction-oriented Prompt Optimization (FIPO). The modular FIPO schema proposes to address Automatic Prompt Optimization as end-to-end text generation, flexibly taking raw task instruction, raw response and ground truth as inputs, for obtaining a new optimized instruction-oriented prompt. We hereby collect a large-scale, high-quality prompt preference dataset, employ with multiple fine-tuning strategies, and then validate the efficacy across five public benchmarks with various downstream generators.
Limitations
On the one hand, due to the current open-source models primarily concentrating on the “small-scale” around 10 billion parameters and otherwise “large-scale” around 70 billion parameters, we lack the opportunity to experiment with medium-sized models (e.g., 40 billion parameters). Based on the conclusions we have drawn, the 10B “small” LLMs are completely incapable of addressing free-from Automatic Prompt Optimization. Therefore, we would like to explore the critical model size threshold for this capability in our future research.
On the other hand, our current research on FIPO does not include optimization of the in-context examples, but rather focuses on prompt optimization for the task instructions. We maintain an open attitude towards related research.
Appendix A HyperParameters and Training Cost
We report hyperparameters and training cost in Table 6. We fine-tune all models on same computational node server equipped with 900G CPU RAM and 8 A100 40G GPUs. We only use single node for all experiments with 13B models, and LORA training of 70B models. As for full training of 70B models, we takes 4 nodes. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find larger learning rate would destroy the training stability of preference optimization. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).
The neural optimizer that we adopted is AdamWLoshchilov and Hutter (2019) and the scheduler employed is WarmupDecayLRGoyal et al. (2017). We incorporate Deepspeed Ren et al. (2021) and Flash Attention Dao et al. (2022) to improve the training efficiency. It worth to mention that IPL approaches takes more than double training time per epoch compared with regular preference optimization approaches, since we increment additional instruction following data and discrimination data, as well as self-rewarding data updates.
13B-Lora 13B-Full 70B-Lora 70B-Full
Nodes 1 1 1 4
Batch 16 8 2 2
Accumulations 1 2 8 8
HyperParams Epoch=3, Seq Len=2048, SFT Lr=2e-5, Else Lr=5e-7, Warmup Ratio=0.1, Beta=0.01, Gen TopP=0.95, Gen Temperature=0.8
Train (1 epoch) 1.5h 2.5h 6h 2.5h
Table 6: Hyperparameters and training cost.
Figure 5: An overview of our dataset diversification step. It is recommended to view the details with colors.
Appendix B Dataset Diversification
We present our complete data diversification plans in Figure 5. We first collect response and prompt preference data as shown in the top. Afterwards, to better minimize the exposure gaps, we carry on the diversification combinations as shown within the table in the left middle of Figure 5. The diversification optionally selects raw instruction response $\bm{\hat{y}^{r}}$, ground truth response $\bm{y}^{r}$ and organization format. We thereby have eight various data format, and we demonstrate with two types in the bottom left and middle parts. The left corner one is an example format of type 3, in which only mandatory raw task instruction $\bm{x}^{r}$ is included. It worth to mention that this format type is also used during the testing phase as no responses are provided at that stage. Another type 6 located in the middle bottom transforms the initial generation format as a multi-choice one, and add a new binary raw instruction response $\bm{\hat{y}^{r}}$. Nevertheless, either type 3 or type 6 has fixed task description attached at the end.
The right side of Figure 5 illustrates specific incremented data for IPL fine-tuning. Equations 4 and  1 denote that only optimizer model $\bm{M_{o}}$ is included in preference optimization and supervised fine-tuning approaches. However, discrimination and instruction answering capacities are required in our iterative self-rewarding training, as shown in equations 7 and  8. Hence, we simply re-use the collected data in the top of Figure 5. 15k raw task instruction $\bm{x^{r}}$ and ground truth response $\bm{y^{r}}$ are paired as additional insturction following data of IPL, shown in the right middle part. In the meantime, we set a new meta prompt for the inner discrimination of IPL, shown in the right corner of Figure 5. Another 15k raw task instruction $\bm{x^{r}}$, ground truth response $\bm{y^{r}}$ and corresponding prompt labels $\bm{x^{o-}}$ and $\bm{x^{o+}}$ are re-used for data construction.
In summary, we diversify the collected 30k raw data into eight various formats, for bridging the gaps between training and inference. Particularly, we add 15k instruction following data and 15k discrimination data for IPL, by re-using the raw data.
1:  Input require: Total number of iterations $\bm{N}$,
2:  Optimizer model $\bm{M_{o}}$,
3:  Discriminator model $\bm{M_{j}}$,
4:  Generator model $\bm{M_{g}}$,
5:  $\bm{M_{o}}\equiv\bm{M_{j}}\equiv\bm{M_{g}}$.
6:  Initilization: Raw Task Instruction $\bm{x^{r}}$,
7:  Raw Task Response $\bm{\hat{y}^{r}}$,
8:  Ground Truth Response $\bm{y}^{r}$,
9:  Suboptimal Optimized Prompt $\bm{x^{o-}}$,
10:  Optimal Optimized Prompt $\bm{x^{o+}}$.
11:  for $\bm{n}$ in $\bm{N}$ do
12:     if $\bm{n}$ > 1 then
13:        Get a prompt $\bm{x^{r+}}=\bm{M_{o}}(\bm{x^{r}},\bm{\hat{y}^{r}},\bm{y}^{r})$
14:        if $\bm{M_{j}}(\bm{x^{r}},\bm{y^{r}})\succeq\bm{M_{j}}(\bm{x^{r+}},\bm{y^{r}})$ then
15:           Reassign $\bm{x^{r+}}=\bm{x^{r}}$
16:        end if
17:        Get new response $\bm{\hat{y}^{r+}}=\bm{M_{g}}(\bm{x^{r+}})$
18:        Update raw task instruction $\bm{x^{r}}=\bm{x^{r+}}$
19:        Update raw task response $\bm{y}^{r}=\bm{\hat{y}^{r+}}$
20:     end if
21:     Update $\bm{M_{o}}$ with DPO or IPO loss.
22:  end for
Algorithm 1 Self-rewarding IPL Algorithm.
Appendix C Additional details of IPL
We first report the algorithmic narrative of IPL in Algorithm 1. In consistent with Equations 5,  6,  7 and  8, we first generate a new prompt $\bm{x^{r+}}$ with the optimizer model $\bm{M_{o}}$, then ask the model itself to judge if the newly optimized prompt is superior than the raw prompt $\bm{x^{r}}$. We takes the better prompt as the final one, and then generate a new task response $\bm{\hat{y}^{r+}}$ for next iteration. It should be noted that we start such self-rewarding updates after one epoch of warm-up.
We separately report representative accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in our IPL training pipeline in Table 7. We takes the fulling training of 70B model with IPO loss as the example, noted as “IPL-IPO-70B-eN (F)”, in which “-eN” refers to the “N”-th round of iteration. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. Nevertheless, the optimizer model itself gradually adds new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. The first epoch is for warm-up as aforementioned, therefore there are not scores reported.
Weighted Avg. Dis. Accuracy Selection
Human 47.80 N/A N/A
IPL-IPO-70B-e1 (F) 48.30 N/A N/A
IPL-IPO-70B-e2 (F) 50.81 100% 1.25%
IPL-IPO-70B-e3 (F) 52.13 100% 2.40%
Table 7: Analysis regards to details of IPL. We takes full training of 70B model with IPO loss as the example. The “-eN” notation refers to the “N”-th round of iteration in IPL training. The selection refers to the data proportion that be updated with newly generated prompts.
Appendix D Entire fine-tuning results
Generation Multi-choice
FIPO Optimizer GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Human 24.77 36.21 73.35 51.17 51.22 47.79
13B (L) 18.62 29.98 71.61 46.95 46.94 43.14
SFT 13B (F) 19.91 31.04 71.61 47.68 49.57 44.93
70B (L) 26.44($\uparrow$) 32.46 73.90($\uparrow$) 50.64 50.03 46.33
70B (F) 21.43 32.92 74.39($\uparrow$) 49.97 51.55($\uparrow$) 46.96
13B (L) 26.98($\uparrow$) 34.62 71.06 47.48 50.53 46.59
DPO 13B (F) 17.17 30.90 68.50 44.87 46.64 42.69
70B (L) 24.24 41.88($\uparrow$) 76.24($\uparrow$) 55.87($\uparrow$) 54.09($\uparrow$) 51.39($\uparrow$)
70B (F) 27.74(normal-$\uparrow$) 35.56 74.17($\uparrow$) 54.93($\uparrow$) 52.73($\uparrow$) 49.07($\uparrow$)
13B-e1 (L) 27.05($\uparrow$) 35.75 70.95 48.99 50.77 47.16
13B-e2 (L) 24.92($\uparrow$) 35.88 71.28 47.65 51.17 47.17
13B-e3 (L) 25.61($\uparrow$) 35.23 70.41 49.56 51.10 47.16
70B-e1 (L) 22.04 42.15($\uparrow$) 75.26($\uparrow$) 55.64($\uparrow$) 55.02($\uparrow$) 51.74($\uparrow$)
IPL-DPO 70B-e2 (L) 23.48 42.63($\uparrow$) 75.97($\uparrow$) 56.11($\uparrow$) 54.09($\uparrow$) 51.54($\uparrow$)
70B-e3 (L) 21.96 42.04($\uparrow$) 75.64($\uparrow$) 56.01($\uparrow$) 55.13($\uparrow$) 51.84($\uparrow$)
70B-e1 (F) 25.38($\uparrow$) 33.86 74.39($\uparrow$) 52.42($\uparrow$) 52.44($\uparrow$) 48.12($\uparrow$)
70B-e2 (F) 24.28 36.10 73.84($\uparrow$) 51.22($\uparrow$) 52.04($\uparrow$) 48.23($\uparrow$)
70B-e3 (F) 25.13($\uparrow$) 35.25 74.95($\uparrow$) 50.46 52.12($\uparrow$) 48.10($\uparrow$)
13B (L) 24.62 34.40 71.72 48.99 50.65 46.70
IPO 70B (L) 22.95 42.50($\uparrow$) 76.95($\uparrow$) 55.97($\uparrow$) 54.17($\uparrow$) 51.58($\uparrow$)
70B (F) 25.00($\uparrow$) 39.21($\uparrow$) 76.84($\uparrow$) 56.01($\uparrow$) 54.29($\uparrow$) 50.94($\uparrow$)
13B-e1 (L) 24.47 35.86 72.15 47.95 50.68 46.98
13B-e2 (L) 26.98($\uparrow$) 35.30 71.93 48.86 50.36 46.88
13B-e3 (L) 25.38($\uparrow$) 35.93 71.23 49.03 51.18 47.36
70B-e1 (L) 22.19 41.88($\uparrow$) 75.53($\uparrow$) 55.81($\uparrow$) 54.12($\uparrow$) 51.25($\uparrow$)
IPL-IPO 70B-e2 (L) 22.11 42.77($\uparrow$) 75.59($\uparrow$) 55.47($\uparrow$) 54.56($\uparrow$) 51.66($\uparrow$)
70B-e3 (L) 21.96 42.92(normal-$\uparrow$) 75.37($\uparrow$) 55.64($\uparrow$) 55.03($\uparrow$) 51.94($\uparrow$)
70B-e1 (F) 25.99($\uparrow$) 34.09 74.66($\uparrow$) 53.62($\uparrow$) 52.33($\uparrow$) 48.30($\uparrow$)
70B-e2 (F) 26.70($\uparrow$) 38.07($\uparrow$) 76.28($\uparrow$) 54.38($\uparrow$) 54.81($\uparrow$) 50.81($\uparrow$)
70B-e3 (F) 26.67($\uparrow$) 39.60($\uparrow$) 77.11(normal-$\uparrow$) 56.71(normal-$\uparrow$) 56.02(normal-$\uparrow$) 52.13(normal-$\uparrow$)
Table 8: Complete Evaluation results of downstream testing on out-of-box Tulu2-7B model. “Human” refers to use raw few-shot prompts written by humans, while other lines are tests with optimized prompts. The number attached with the benchmark name refers to the number of in-context examples (e.g., BBH (3) stands for 3-shot evaluation on BBH). The “(L)” or “(F)” letter attached with FIPO optimizer refer to parameter-efficient tuning LORA Hu et al. (2022) and full-parameter tuning, respectively. The “X” part of “IPL-X” refers to the different loss style mentioned in Equation 5. And the “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model).
We present entire fine-tuning results of our various experiments in Table 8. Apart from conclusions reported in Section 4.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected, so as to the full training compared with Lora. However, all SFT models perform worse than the human baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe the Lora tuning performs better than the full-parameter training on the contrary. Nevertheless, 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO, as we find Lora tuning with 70B models behave the best, although our IPL strategy can improve with margins. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend. We suppose the regularized design in Equation 3 contributes to stable full-parameter improvements.
Appendix E Employed Meta Prompts
We report all employed meta prompts in this section as reference for potential future researches. Please refer to Table 9,  10 and  11 for data collection meta prompt, data diversification meta prompt, and discrimination meta prompt, respectively.
Data collection meta prompt
You are an expert of prompt optimization.
Sliver Prompt:
SP
Sliver Response:
SR
Golden Response:
GR
Task Introduction:
Based on the Silver Prompt, optional Silver Response and optional Golden Response, perform the following actions:
1 – The optional Sliver Response was generated by an AI based on the Silver Prompt. Please help modify the Silver Prompt to Golden Prompt that can obtain a more correct response, in reference to the optional Golden Response.
2 - When building the Golden Prompt, you can consider several aspects, such as: (1) A roleplay leading sentence to adapt the AI to the task-specific scenario; (2) Details of task characteristics, for instance, the task could be a question answering task, a dialogue task, or a summarization task, etc; (3) Further clarification of the task information, especially some ambiguous terms; (4) A more detailed solution guidance, such as step-by-step plans, handlings of exceptions, special priorities or constraints, etc; (5) Any specific requirements for the response, such as the length, the format, the style, the tone, the language, etc.
3 - Show me only the Golden Prompt, do not contain any other content.
Golden Prompt:
Table 9: The meta prompt used to harness data from GPT-3.5 and GPT-4 APIs. During harnessing, we replace placeholders “SP”, “SR” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Response $\bm{\hat{y}^{r}}$ and Ground Truth Response $\bm{y}^{r}$, respectively.
Data diversification meta prompt
You are an expert of prompt optimization.
Sliver Prompt:
SP
<Optional Responses>
The optional Sliver Response was generated by an AI based on the Silver Prompt. Please help modify the Silver Prompt to Golden Prompt that can obtain a more correct response, in reference to the optional Golden Response. The Golden Prompt should be instructive, concise and strictly faithful to any factual information in the Silver Prompt. The length of the Golden Prompt should be less than GN words. Only give me the content of Golden Prompt, do not contain any other information (e.g., your response of the Golden Prompt, any postfix like “Golden Prompt”, etc.).
Flexible meta prompt for optional raw task response
Sliver Response:
SR
Flexible meta prompt for optional ground truth response
Golden Response:
GR
Table 10: The meta prompt used for data diversification. During diversification, we replace placeholders “SP”, “SR” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Response $\bm{\hat{y}^{r}}$ and Ground Truth Response $\bm{y}^{r}$, respectively. Meanwhile, flexible prompts of optional raw task response and ground truth response are inserted by probability as introduced in Figure 3 and  5. Moreover, we add length suggestion with “GN” signal, using the length of optimal optimized prompt label $\bm{x^{o+}}$.
Discrimination meta prompt
You are an expert of prompt discrimination.
Raw Prompt:
RP
New Prompt A:
PA
New Prompt B:
PB
Golden Response:
GR
New Prompt A and New Prompt B are optimized from Raw Prompt. Please judge which prompt is more loyal to the factual information of Raw Prompt, and is more desirable for an AI to generate the Golden Response. Only answer with A or B.
Table 11: The meta prompt used for creating discrimination data and application of discrimination during IPL training. The creation of discrimination data has been introduced in the right part of Figure 5. In terms of inner discrimination of IPL, we replace placeholders “RP”, “PA”, “PB” and “GR” with actual Raw Task Instruction $\bm{x^{r}}$, Raw Task Instruction $\bm{x^{r}}$, Newly Optimized Prompt $\bm{x^{r+}}$ and Ground Truth Response $\bm{y}^{r}$, respectively.