### v3

# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

Abstract
When the quality of naive prompts is carefully optimized by human experts, the task performance of large language models (LLMs) can be significantly improved. However, expert-based prompt optimizations are expensive. Herein, some works have proposed Automatic Prompt Optimization (APO), to optimize naive prompts according to task outputs of given in-box testing models, with the help of advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, existing schemes suffer from poor generalization ability and privacy risk. To this end, we collect the first large-scale Prompt Optimization Preference dataset (POP), fine-tune offline local LLM-based optimizers, then fairly test with various downstream models. Our method allows accurate optimization of the core task instruction part within the naive prompt in a model-agnostic manner, and thus is named Free-from Instruction-oriented Prompt Optimization (FIPO). In specific, FIPO uses a modular APO template that dynamically integrate the naive task instruction, optional instruction responses, and optional ground truth to produce finely optimized prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. Leveraging insights from the data with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of FIPO framework across five public benchmarks and six testing models [^1] Check codes and data here: https://github.com/LuJunru/FIPO_Project..
FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema
Junru Lu1, Siyu An2, Min Zhang3, Yulan He4, Di Yin2, Xing Sun2 1University of Warwick, 2Tecent YouTu Lab, 3Westlake University, 4King’s College London 1junru.lu@warwick.ac.uk, 3zhangmin@westlake.edu.cn, 4yulan.he@kcl.ac.uk 2{siyuan, endymecyyin, winfredsun}@tencent.com
1 Introduction
Large Language Models (LLMs) show their intelligence Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) on various benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). However, the task performance of LLMs depends on the given task prompt. LLMs may not be able to give the correct answer based on the naive task prompt, but after careful optimization by human experts, it can perform the same task well under the guidance of high-quality prompts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b).
Figure 1: Online ad-hoc APO vs. our Local End2end FIPO. Although both leverage advanced LLMs (e.g., GPT-4), in FIPO, we design a local training pipeline and completely removed the reliance on any in-box model generators.
Obviously, expert-based prompt optimizations are expensive. Therefore, in recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, focusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce ad-hoc strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)).
These APO approaches typically involve iterative optimization between an in-box testing generator $\bm{M}_{g-in}$ and an advanced optimizer $\bm{M}_{o-api}$. As illustrated in the upper half of Figure 1, the generator $\bm{M}_{g-in}$ first provide answers upon the naive prompt $\bm{x}^{n}$: “Calculate the average value of the list”, and then the optimizer $\bm{M}_{o-api}$ provides rational feedback and several upgraded prompt candidates $\{\bm{x}^{o}\}$. This above process can be repeated several times. Afterwards, a high-quality optimized prompt $\bm{x}^{o}$ is finally tailored for the in-box generator $\bm{M}_{g-in}$, which can be ensured to produce the desired response: “The answer value is 44.25”. However, despite effective, several drawbacks exist: (1) Privacy risk. The entire online optimization is exposed to external LLM services, which poses a privacy risk; (2) Poor Generalization Ability. The ad-hoc optimization is model-centric, as it runs with instant testing responses from the in-box generator $\bm{M}_{g-in}$, leading to performance degradation when tested with out-of-box generators $\bm{M}_{g-out}$. For instance, a wrong answer is inferred by the out-of-box generator $\bm{M}_{g-out}$: “44.1”.
To address the existing challenges, we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). The bottom half of Figure 1 illustrates our FIPO. Unlike the online ad-hoc APO approach, FIPO directly fine-tunes the general local optimizer $\bm{M}_{local}$, and apply over any testing generator $\bm{M}_{g}$. In specific, we first design a meta template towards universal APO (Figure 2), and thereby collect 30,000 prompt optimization preference examples from advanced optimizer $\bm{M}_{o-api}$ (§ 2.3). We demonstrate the reliability of this data through multiple cross-validation methods (Table 1). On this basis, we then apply with mainstream end-to-end fine-tuning strategies for obtaining a valid local optimizer $\bm{M}_{o-local}$ (Figure 3). Particularly, in the Figure 1’s example, our local optimizer $\bm{M}_{local}$ provides insightful hints in the optimized prompt $\bm{x}^{o}$: “calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list”. Such stepwise guidance generally enhance the answer quality of any testing generator $\bm{M}_{g}$. Our contributions are as follows:
  • (1)
    We denote the drawbacks of previous online ad-hoc APO, and then introduce our local free-form instruction-oriented prompt optimization (FIPO). We design a generic APO template that allows the optimizer to unpack the pre-built testing generators, which is a key component for former methods.
  • (2)
    Based on the template, we compile a 30k large-scale Prompt Optimization Preference dataset (POP), and explore with several mainstream fine-tuning strategies: Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). We further develop an Iterative Preference Learning (IPL) pipeline for self-enhanced preference optimization.
  • (3)
    We evaluate the efficacy and adaptability of fine-tuned FIPO optimizer across five downstream benchmarks and three diverse generators, and obtain superior results compared to previous online ad-hoc APO methods.
Figure 2: Step 1 and 2 of FIPO: (1) Design a meta template towards universal APO; (2) Collect 30,000 large-scale prompt optimization preference exemplars from a suboptimal LLM GPT-3.5-turbo and an optimal LLM GPT-4.
2 Methodology
In this section, we start with task formulation (§ 2.1), then introduce our meta template towards universal APO (§ 2.2), collected POP data (§ 2.3) and training strategies (§ 2.4).
2.1 Task Formulation
We denote FIPO as end-to-end text generation. In the training phase, a local optimizer model $\bm{M}_{o-local}$ is supervisedly fine-tuned to generate an optimized prompt $\bm{\hat{x}}^{o}$:
$\bm{\hat{x}}^{o}=\operatorname*{argmax}_{\bm{M}_{o-local}}p(\bm{\hat{x}}^{o}|\bm{x}^{n},[\bm{\hat{y}}^{n}],[\bm{y}^{n}])$ (1)
based on the naive prompt $\bm{x}^{n}$, optional naive response $\bm{\hat{y}}^{n}$, and optional ground truth $\bm{y}^{n}$. In addition, pairwise chosen optimized prompt $\bm{x}^{o+}$ and rejected optimized prompt $\bm{x}^{o-}$ are provided as labels in training. The optional naive response $\bm{\hat{y}}^{n}$ is generated for the naive prompt $\bm{x}^{n}$ using one neural generator model $\bm{M}_{g*}$:
$\bm{\hat{y}}^{n}=\operatorname*{argmax}_{\bm{M}_{g*}}p(\bm{\hat{y}}^{n}|\bm{x}^{n})$ (2)
While in the testing phase, our ultimate target is to obtain a more superior optimized testing response $\bm{\hat{y}}^{o}_{t}$ than the naive testing response $\bm{\hat{y}}^{n}_{t}$, when applying any testing generator $\bm{M}_{g}$ to the optimized testing prompt $\bm{\hat{x}}^{o}_{t}$ and the naive testing prompt $\bm{x}^{n}_{t}$, respectively:
$\bm{\hat{y}}^{o}_{t}\succ\bm{\hat{y}}^{n}_{t}$ (3)
$\bm{\hat{y}}^{o}_{t}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{o}_{t}|\bm{\hat{x}}^{o}_{t})\mbox{, }\bm{\hat{y}}^{n}_{t}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{n}_{t}|\bm{\hat{x}}^{n}_{t})$ (4)
where $\bm{M}_{g}$ could be either same as or different from $\bm{M}_{g*}$. And specifically, $\bm{\hat{x}}^{o}_{t}$ is enhanced from $\bm{x}^{n}_{t}$ by the fine-tuned optimizer $\bm{M}_{o-local}$:
$\bm{\hat{x}^{o}_{t}}=\operatorname*{argmax}_{\bm{M_{o}}}p(\bm{\hat{x}}^{o}_{t}|\bm{x}^{n}_{t})$ (5)
In contrast, former ad-hoc APO has no training phase but only the iterative online testing pipeline with mandatory in-box testing response $\bm{\hat{y}}^{o_{i}}_{t}$:
$\bm{\hat{x}}^{o_{i+1}}_{t}=\operatorname*{argmax}_{\bm{M}_{o-api}}p(\bm{\hat{x}}^{o_{i+1}}_{t}|\bm{x}^{o_{i}}_{t},\bm{\hat{y}}^{o_{i}}_{t})\mbox{, }\bm{x}^{o_{1}}_{t}=\bm{x}^{n}_{t}$ (6)
$\bm{\hat{y}}^{o_{i}}_{t}=\operatorname*{argmax}_{\bm{M}_{g-in}}p(\bm{\hat{y}}^{o_{i}}_{t}|\bm{x}^{o_{i}}_{t})\mbox{, }\bm{\hat{y}}^{o_{1}}_{t}=\bm{\hat{y}}^{n}_{t}$ (7)
where $\bm{M}_{g-in}$ is a prior in-box generator.
2.2 Modular Template
As aforementioned in section 2.1, we first design a modular template that ensures flexibility in content management. Figure 2 illustrates our template, shown in the middle, taking mandatory naive task instruction $\bm{x}^{n}$, optional naive response $\bm{\hat{y}}^{n}$ and optional ground truth response $\bm{y}^{n}$ as inputs. Additional description is then appended for clarity, in which we directly claim the optionality of $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$: “The optional sliver response ··· based on the sliver prompt ··· optional golden response ···”.
We use this modular template for all sections in FIPO, including the dataset collection, fine-tuning the local optimizer $\bm{M}_{o-local}$, and testing various downstream generators $\{\bm{M}_{g}\}$. The key difference between these sections is we accordingly adjust the optional responses, thus addressing potential exposure bias between the training and the inference phases (Eq. 1 vs. Eq. 5): (1) Keep both in data collection. We introduce the collection of our POP data in section 2.3; (2) Diversely keep partial responses in training. We present this strategy in section 2.4; (3) Remove all responses in testing. We report testing results in section 3.2.
Figure 3: Step 3 of FIPO: transitional dataset diversification and several mainstream fine-tuning strategies.
2.3 Prompt Optimization Preference Data
We decide to distill the refined prompt optimization capabilities from prominent yet proprietary LLMs, instead of directly integrating them in an ad-hoc way. Thus, we collect the Prompt Optimization Preference (POP) data. Shown in Figure 2, to ensure the most directional optimization, we send naive prompt $\bm{x}^{n}$, naive response $\bm{\hat{y}}^{n}$, and ground truth response $\bm{y}^{n}$ to one suboptiomal LLM GPT-3.5-turbo and one optimal LLM GPT-4222https://platform.openai.com/docs/models, to collect contrastive POP data ($\bm{x}^{o+}$, $\bm{x}^{o-}$). The naive prompt $\bm{x}^{n}$ is sampled from the Alpaca dataset, which contains 52K diverse instructions and corresponding responses $\bm{\hat{y}}^{n}$ generated by the Text-davinci-003 model Taori et al. (2023). We also collect another GPT4-generated response for the Alpaca dataset from public literature Ghosal et al. (2023). There are no official ground truth responses (e.g., from human experts) for the Alpaca data, we therefore consider GPT-4 responses as the ground truth response $\bm{y}^{n}$, given its demonstrated analytical capabilities comparable to humans Pan et al. (2023a). As shown at the bottom of Figure 2, GPT-4 offers a more pedagogical step-by-step optimized prompt compared to GPT-3.5-turbo. We report complete collection template in Table 10.
We finally narrow down our dataset to 30k samples and report the quality post-checking in Table 1. We adopt cross-validation using three different methods: critiques from an external alignment model UltraRM Cui et al. (2023), self-judgement from the GPT-4, and manual checking by human experts. The “Response” and “Prompt” columns refer to the proportions that GPT-4-generated response and GPT-4-optimized prompt is better than the others, respectively. The average win rates for both categories exceed 85%, ensuring the quality.
GPT-4 Win Rate (%)
Response Prompt Scale
UltraRM 13B Cui et al. (2023) 91.49 82.13 30k
GPT4 Self-check 80.56 92.29 3k
Human Expert 88.29 95.21 1k
Average 86.78 89.88 N/A
Table 1: Quality cross-validation on our dataset.
2.4 Fine-tuning Strategies
We introduce our fine-tuning strategies in Figure 3, consisting of an initial step of transitional dataset diversification followed by strategic fine-tuning.
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix B..
Strategic Fine-tuning. The right side of Figure 3 introduces several end-to-end fine-tuning strategies that we explored in this work. The right top is the most well-known Supervised Fine-tuning (SFT), which only takes the optimal optimized prompt $\bm{x}^{o+}$ as the supervision signal:
$L_{SFT}(M_{o})=-\mathbb{E}_{(x^{n},\hat{y}^{n},y^{n},x^{o+})\sim D}[\hat{x}^{o}-x^{o+}]^{2}$ (8)
where $D$ stands for the training set.
On the other hand, the right middle shows a contrastive fine-tuning methodology: Preference Optimization, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). Preference Optimization takes pairwise rejected label $\bm{x}^{o-}$ and chosen label $\bm{x}^{o+}$ as supervision. One of the core differences from Preference Optimization and SFT is the former one not only encourage the generation of optimal preference, but also dampen the generation of suboptimal preference:
$L_{DPO}(M_{o})=-\mathbb{E}_{(x^{n},\hat{y}^{n},y^{n},x^{o+},x^{o-})\sim D}[\log\sigma(\beta\cdot\Delta)]$ (9)
$L_{IPO}(M_{o})=-\mathbb{E}_{(x^{n},\hat{y}^{n},y^{n},x^{o+},x^{o-})\sim D}[\Delta-\frac{1}{2\beta}]^{2}$ (10)
$\Delta=\log\frac{M_{o}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}-\log\frac{M_{o}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}$ (11)
where $\bm{\beta}$ is a hyperparameter factor. $\bm{M_{ref}}$ refers to the reference model, which is a frozen copy of initial weights of $\bm{M_{o}}$. The equations indicate that IPO is a regularized version of DPO as it limits the optimization range with squares.
Additionally, inspired by self-updating alignment Lee et al. (2023); Anthropic (2022); Yuan et al. (2024), we develop a Iterative Preference Learning (IPL) strategy for self-rewarding prompt optimization. After each iteration of prompt optimization, we ask the optimizer itself to determine if it successfully generate a superior prompt $\bm{x}^{n+}$ with a better response $\bm{\hat{y}}^{n+}$, and if so, to automatically replace the previous inferior prompt $\bm{x}^{n}$ and response $\bm{\hat{y}}^{n}$, leading to more rigorous training in next iteration:
$L_{IPL}(M_{o})=-\mathbb{E}_{(x^{n+},\hat{y}^{n+},y^{n},x^{o+},x^{o-})\sim D}G(\Delta)$ (12)
$x^{n+}=M_{o}(x^{n})\mbox{, }\hat{y}^{n+}=M_{o}(x^{n+})$ (13)
$x^{n+}=\left\{\begin{array}[]{l}x^{n+}\mbox{, }M_{o}(x^{n+},y^{n})\succ M_{o}(x^{n},y^{n})\\ x^{n}\mbox{, otherwise}\end{array}\right.$ (14)
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix B and  C..
3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix A..
3.1 Experimental Settings
3.1.1 Baselines
We compare FIPO with two SOTA APO methods: APE Zhou et al. (2023) and PromptAgent Wang et al. (2023a). APE, stands for Automatic Prompt Engineer, is a template-based strategy that ask one LLM to generate a pool of candidate prompts based on the templates, then select one according to the evaluation scores. PromptAgent eliminates templates and replaces with Monte Carlo Tree Search Abramson (2014) for using a evaluator model to guide the generator. Both APE and PromptAgent are training-free, aiming to model-oriented APO in an ad-hoc manner, while we realize completely offline training. Following former works, we use GPT-3.5-turbo as the in-box generator and GPT-4 as the optimizer in both baselines.
3.1.2 Foundation models
We use Tulu2 models as our bases, which is a fine-tuned version of Llama2 Touvron et al. (2023) trained on a mix of publicly available datasets Ivison et al. (2023). We fine-tuning local optimizer with Tulu2-13B and Tulu2-70B.
Generation Multi-choice
Generator Prompt Source GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Llama2-7B Naive Prompt 8.89 31.21 62.78 43.09 46.58 41.73
Touvron et al. (2023) FIPO Optimizer 11.70 33.50 69.37 52.11 54.56 48.10
Tulu2-13B Naive Prompt 39.06 36.49 76.62 55.13 57.43 52.53
Ivison et al. (2023) FIPO Optimizer 40.17 40.26 78.58 57.68 59.10 54.79
Baichuan2-13B Naive Prompt 46.81 37.95 68.56 51.88 57.46 52.36
Yang et al. (2023a) FIPO Optimizer 48.12 39.95 74.77 56.88 58.32 54.35
Table 2: Evaluation results of various downstream generator LLMs, using the best local optimizer from Table 3.
Figure 4: The FIPO-optimized prompts help various downstream testing LLMs (X-axis) gain more promising improvements, compared with other prompt optimization approaches (shown by the bars). We specifically annotate the improvements of FIPO against naive prompts from the original dataset (). More details: Appendix G and  D.
3.1.3 Evaluation Benchmarks
We include five benchmarks across two most common formats: (1) GSM8k Cobbe et al. (2021), a generative dataset contains 1.3k primary level math questions; (2) BigBenchHard (BBH) Suzgun et al. (2023), which involves 23 challenging reasoning tasks. BBH has 6.4k testing samples, and asks for generative answering; (3) PiQA Bisk et al. (2020), in which 1.8k common physical knowledge questions are proposed, alongside with multiple candidate choices; (4) CosmosQA Huang et al. (2019). There are around 3k commonsense-based multi-choice questions in CosmosQA, equipped with four candidate options; (5) MMLU Hendrycks et al. (2021), which is one of the largest multi-choice benchmarks. MMLU covers 14k questions. For our FIPO experiments, we report results on all five benchmarks. While differently, since both APE and PromptAgent only provide evaluations on 6 tasks of BBH, we report the comparison results aligned with their settings. As for the result metrics, either “generation” or “multi-choice” benchmarks takes few-shot format with strict answering templates (e.g., “The answer is X”). Herein, we are able to report the accuracy score for all benchmarks.
3.2 Experimental Results
3.2.1 Efficacy of FIPO
General improvements of FIPO. It can be concluded that FIPO-optimized prompts have general gains on different downstream generators across five public benchmarks, shown in Table 2. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average.
Comparable optimization capability against online ad-hoc APO, even better. We would like to compare the local FIPO optimizer with previous methods and direct prompt optimization using GPT-4. Figure 4 reports experimental results on six BBH tasks, following the experimental settings in PromptAgent work Wang et al. (2023a). Our FIPO method takes the lead in all downstream tests, except for the in-box tester GPT-3.5-turbo, which is ad-hocly included during the iterative prompt optimization in APE and PromptAgent. In specific, the final average improvements on two open-source 70B models are around 3% to 5%, compared with more than 10% gains on two open-source 7B models. We can notice that as the tested open-soruce model grows larger and stronger, the effectiveness of all prompt optimization methods significantly decreases, which may be due to the firmness of the larger model’s inherent knowledge. As for proprietary GPT-3.5 and GPT-4, we found that prompt optimization seems to be more beneficial to them. Prompts optimized by our FIPO can help GPT3.5 improve the final average effect up to 22%, and maintain an effect of about 8% on GPT4o.
Generation Multi-choice
FIPO Prompt Optimizer GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Naive Prompt 24.77 36.21 73.35 51.17 51.22 47.79
Best 13B Optimizer 18.42 33.55 72.03 49.14 48.20 44.93
SFT-70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO-70B 27.74 35.56 74.17 54.93 52.73 49.07
IPO-70B 25.00 39.21 76.84 56.01 54.29 50.94
IPL-DPO-70B 25.13 35.25 74.95 50.46 52.12 48.10
IPL-IPO-70B 26.67 39.60 77.11 56.71 56.02 52.13
IPO-70B-gen 22.72 41.91 74.53 53.60 52.74 50.23
IPO-70B-partial 23.08 40.03 76.22 54.29 51.99 49.59
Table 3: Evaluation results of FIPO optimizer fine-tuned with different strategies, tested with same Tulu2-7B model. The number attached with the benchmark is the number of in-context examples (e.g., BBH (3) means 3-shot testing on BBH). FIPO only optimizes the last task instruction, leaving in-context examples remained.
3.2.2 Fine-tuning Analysis
We fine-tune Tulu2-13B and -70B models as our local optimizer through different strategies as mentioned above. We use downstream performance of Tulu2-7B for analyzing the effectiveness of different fine-tuning strategies. Our findings are here:
Small optimizer fails. “Small” Tulu2-13B are not up to the difficult prompt optimization task (the 4th line of Table 3). The average testing scores of using optimized prompts even worse than using the naive prompts written by human.
SFT vs. DPO/IPO vs. IPL. When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task. While when contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%. In terms of different preference fine-tuning methods, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline, which may due to its regularized design in Eq. 10. We analyze more fine-tuning details and the self-rewarding benefits of IPL in Appendix D.
Dataset diversification is necessary. In the bottom of Table 3, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The ablated optimizer weakens all benchmarks, except BBH, which is due to its unique symbolic reasoning pattern (Appendix E).
Naive Testing Prompt I $\bm{x}^{n}_{t_{1}}$ from BBH:
Jane was born on the last day of February in 2000. Today is her 16-year-old birthday. What is the date a month ago in MM/DD/YYYY?
Optimized Testing Prompt I $\bm{x}^{o}_{t_{1}}$ by FIPO
As today marks Jane’s 16th birthday, determine the date from exactly one month prior. Jane was born on the last day of February in the year 2000, a leap year. Ensure your response is in the format MM/DD/YYYY.
Naive Testing Prompt II $\bm{x}^{n}_{t_{2}}$ from MMLU:
Three of the following are classic security properties; which one is not? A. Confidentiality, B. Availability, C. Correctness, D. Integrity.
Optimized Testing Prompt II $\bm{x}^{o}_{t_{2}}$ by FIPO
Which of the following is NOT one of the core security principles? A. Confidentiality, B. Availability, C. Correctness, D. Integrity. Provide the letter corresponding to the incorrect option.
Naive Testing Prompt III $\bm{x}^{n}_{t_{3}}$ from CosmosQA:
A team of engineers constructing signal lights for airplanes that they can use to guide them to runways are attempting to determine the brightness needed for the pilot to be able to detect the tower at 1 mile away. They set the light to a test brightness and establish communication with an inbound pilot. When the pilot is 1 mile away from the tower, he says he cannot see the light. In terms of Signal Detection Theory, what would this be called?
Optimized Testing Prompt III $\bm{x}^{o}_{t_{3}}$ by FIPO
According to Signal Detection Theory, when a pilot cannot detect a signal at a set brightness level, it is called a ’miss’. In contrast, it is called a ’hit’. A team of engineers is testing the brightness of signal lights for airplanes to guide them to runways. They establish communication with an inbound pilot and set the light to a test brightness. When the pilot is one mile away from the tower, he reports that he cannot see the light. What term from Signal Detection Theory describes this situation?
Naive Testing Prompt IV $\bm{x}^{n}_{t_{4}}$ from GSM8K:
Lily has a list of numbers: [12, 34, 56, 75]. Calculate the average value of the list.
Optimized Testing Prompt IV $\bm{x}^{o}_{t_{4}}$ by FIPO
To find the average of the given list of numbers, first, you need to add all the numbers in the list, which are: 12, 34, 56, and 75. Add them up, and you will get a sum of 177. Then, divide the sum by the total number of items in the list, which is 4. So, 177 divided by 4 equals 44.25. Therefore, the average of the list [12, 34, 56, 75] is 44.25.
Table 4: Examples from testing benchmarks.
3.2.3 Case Analysis.
In Table 4, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand and apply the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer. We manually confirm that this issue mostly occurs in GSM8K and BBH’s math calculation questions, happening <10%, in Appendix H.
4 Related Work
Automatic Prompt Optimization (APO) is an simple yet effective technique for grabbing potentials of LLMs in various downstream scenarios. Most APO methodologies can be categorized into two types: discrete APO and continuous APO Liu et al. (2023). Discrete APO searches optimized prompts with optimal combinations of discrete tokens Wallace et al. (2019); Shin et al. (2020); Ben-David et al. (2022); Davison et al. (2019); Deng et al. (2022); Zhang et al. (2023); Xu et al. (2022). For instance,  van de Kar et al. (2022) employed text mining for searching candidate prompts from knowledge triplets. While  Yuan et al. (2021),  Haviv et al. (2021) and  Gao et al. (2021) utilized Bart Lewis et al. (2019), Bert Devlin et al. (2018) and T5 Raffel et al. (2019) for optimizing prompts in an paraphrasing manner, respectively.
In contrast, continuous APO proposes to search better prompts in continuous embedding space rather than limited to human-understandable discrete tokens Tsimpoukelli et al. (2021); Zhong et al. (2021); Qin and Eisner (2021); Hambardzumyan et al. (2021); Wang et al. (2023b). Prefix Tuning Li and Liang (2021) and Prompt Tuning Lester et al. (2021) are two well-known continuous APO methods that insert prefix vectors in front of the task sequence, and then update their corresponding parameters. There are also hybrid works to insert continuous embedding into discrete templates Liu et al. (2021); Han et al. (2021).
Preference Optimization for LLMs allows LLMs to align with human minds in a more nuanced way Pan et al. (2023b), compared to SFT. Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). A series of following work of DPO are further presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024).
5 Conclusion
We introduce FIPO, the Free-form Instruction-oriented Prompt Optimization. The modular FIPO template proposes to address APO as end-to-end text generation, flexibly taking naive prompt, naive response and ground truth as inputs, for obtaining a new optimized prompt. We hereby collect a large-scale prompt optimization preference dataset, employ with multiple fine-tuning strategies, and then validate the efficacy across objective benchmarks with various downstream generators.
Limitations
While FIPO demonstrates significant potential in optimizing prompts for various downstream tasks, there are several limitations to consider:
  • Overwhelmed Cheating Notes. As observed in the case analysis, FIPO sometimes provides overly detailed instructions that can be considered as “cheating notes”. This issue is particularly prevalent in tasks involving mathematical calculations. While this enhances performance, it may not align with the intended use of prompt optimization.
  • Evaluation Metrics. The current evaluation primarily focuses on accuracy metrics. While accuracy is important, other aspects such as the interpretability, fairness, and ethical implications of the optimized prompts should also be considered in future work.
  • Optimization of In-context Examplars. Our current research on FIPO does not include optimization of the in-context examples, but rather focuses on prompt optimization for the task instructions.
Appendix A HyperParameters and Training Cost
We report hyperparameters and training cost in Table 5. We fine-tune all models on same computational node server equipped with 900G CPU RAM and 8 X A100 40G GPUs. We only use single node for fine-tuning 13B models, but use 4 nodes for 70B models. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find large learning rate for DPO and IPO will cause collapse. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).
The neural optimizer that we adopted is AdamWLoshchilov and Hutter (2019) and the scheduler employed is WarmupDecayLRGoyal et al. (2017). We incorporate Deepspeed Ren et al. (2021) and Flash Attention Dao et al. (2022) to improve the training efficiency. It worth to mention that IPL approaches takes more than double training time per epoch compared with regular preference optimization approaches, since we increment additional instruction following data and discrimination data, as well as self-rewarding data updates.
Tulu2 13B Tulu2 70B
Nodes 1 4
Batch 8 2
Accumulations 2 8
HyperParams Epoch=3, Seq Len=2048, SFT Lr=2e-5, Else Lr=5e-7, Warmup Ratio=0.1, Beta=0.01, Gen TopP=0.95, Gen Temperature=0.8
Train (1 epoch / 3W) 2.5h 2.5h
Table 5: Hyperparameters and training cost.
Figure 5: An overview of our dataset diversification step. It is recommended to view the details with colors.
Appendix B Dataset Diversification
We present our complete data diversification plans in Figure 5. We first collect response and prompt preference data as shown in the top, according to pipeline introduced in section 2.4. Afterwards, to better minimize the exposure gaps, we carry on the diversification combinations as shown within the table in the left middle of Figure 5. The diversification optionally selects naive response $\bm{\hat{y}^{n}}$, ground truth response $\bm{y}^{n}$ and “generation” or “multi-choice” format. We thereby have eight various data format ($2\times 2\times 2$), and we demonstrate with two types in the bottom left and middle parts. The left corner one is an example format of type 3, in which only mandatory naive prompt $\bm{x}^{n}$ is included. It worth to mention that this format type is also used during the testing phase as no responses are provided at that stage. Another type 6 located in the middle bottom transforms the initial generation format as a multi-choice one, and add a new binary naive response $\bm{\hat{y}^{n}}$. Nevertheless, either type 3 or type 6 has same task description attached.
The right side of Figure 5 illustrates specific incremented data for IPL fine-tuning. Equations 11 and  8 denote that only optimizer model $\bm{M_{o}}$ is included in preference optimization and supervised fine-tuning approaches. However, discrimination and instruction answering capacities are required in our iterative self-rewarding training, as shown in equations 14. Hence, we simply re-use the collected data in the top of Figure 5. 15k naive prompts $\bm{x^{n}}$ and ground truth responses $\bm{y^{n}}$ are paired as additional instruction following data of IPL, shown in the right middle part. Meanwhile, we set a new meta prompt for the inner discrimination of IPL, shown in the right corner of Figure 5. Another 15k naive prompts $\bm{x^{n}}$, ground truth responses $\bm{y^{n}}$ and corresponding optimized prompts $\bm{x^{o-}}$ and $\bm{x^{o+}}$ are re-used for data construction.
In summary, we diversify the collected 30k raw data into eight various formats, for bridging the gaps between training and inference. Particularly, we add 15k instruction following data and 15k discrimination data for IPL, by re-using the raw data.
1:  Input require: Total number of iterations $\bm{E}$,
2:  Optimizer model $\bm{M_{o}}$,
3:  Initilization: Naive Prompt $\bm{x^{n}}$,
4:  Naive Response $\bm{\hat{y}^{n}}$,
5:  Ground Truth Response $\bm{y}^{n}$,
6:  for $\bm{e}$ in $\bm{E}$ do
7:     if $\bm{e}$ > 1 then
8:        New prompt $\bm{x^{n+}}=\bm{M_{o}}(\bm{x^{n}},\bm{\hat{y}^{n}},\bm{y}^{n})$
9:        if $\bm{M_{o}}(\bm{x^{n+}},\bm{y^{n}})\succ\bm{M_{o}}(\bm{x^{n}},\bm{y^{n}})$ then
10:           New response $\bm{\hat{y}^{n+}}=\bm{M_{o}}(\bm{x^{n+}})$
11:           Update $\bm{x^{n}}=\bm{x^{n+}}$
12:           Update $\bm{\hat{y}^{n}}=\bm{\hat{y}^{n+}}$
13:        end if
14:     end if
15:     Update $\bm{M_{o}}$ with DPO or IPO loss.
16:  end for
Algorithm 1 Self-rewarding IPL Algorithm.
Generation Multi-choice
FIPO Optimizer GSM8K (3) BBH (3) PiQA (3) CosmosQA (5) MMLU (5) Weighted Avg.
Naive 24.77 36.21 73.35 51.17 51.22 47.79
SFT 13B 19.91 31.04 71.61 47.68 49.57 44.93
70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO 13B 17.17 30.90 68.50 44.87 46.64 42.69
70B 27.74 35.56 74.17 54.93 52.73 49.07
70B-e1 25.38 33.86 74.39 52.42 52.44 48.12
IPL-DPO 70B-e2 24.28 36.10 73.84 51.22 52.04 48.23
70B-e3 25.13 35.25 74.95 50.46 52.12 48.10
IPO 13B 18.42 33.55 72.03 49.14 48.20 44.93
70B 25.00 39.21 76.84 56.01 54.29 50.94
70B-e1 25.99 34.09 74.66 53.62 52.33 48.30
IPL-IPO 70B-e2 26.70 38.07 76.28 54.38 54.81 50.81
70B-e3 26.67 39.60 77.11 56.71 56.02 52.13
IPL-IPO-gen 70B-e3 22.72 41.91 74.53 53.60 52.74 50.23
IPL-IPO-partial 70B-e3 23.08 40.03 76.22 54.29 51.99 49.59
Table 6: Complete evaluation results of different fine-tuning strategies, tested by the same Tulu2-7B model. The “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model). “-gen” and “-partial” refer to two ablation experiments in accordance with Section 3.2.2.
Appendix C Additional details of IPL
We first report the algorithmic narrative of IPL in Algorithm 1. In consistent with Equations 12,  13, and  14, we first generate a new prompt $\bm{x^{n+}}$ with the optimizer model $\bm{M_{o}}$, then ask the optimizer also to judge if the newly optimized prompt is superior than the naive prompt $\bm{x^{n}}$. We takes the better prompt as the final one, and then generate a new task response $\bm{\hat{y}^{n+}}$ using the optimizer itself for next iteration. It should be noted that we start such self-rewarding updates after one epoch of warm-up.
we report the accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in IPL in Table 7. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. We use 5% training data as the validation set, and find 100% classification accuracy. Based on this, the optimizer accurately update naive prompts $\bm{x}^{n}$ with self-generate new prompts $\bm{x}^{n+}$, introduced from line 9 to line 11 in Algorithm 1. Nevertheless, the optimizer model gradually update new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. We observe only 2.4% naive prompts and responses samples are upgraded in the end, while the self-game approach significantly promotes the optimization of downstream test prompts, as mentioned in Table 6.
Weighted Avg. Dis. Accuracy Selection
Naive 47.80 N/A N/A
IPL-IPO-70B-e1 48.30 N/A N/A
IPL-IPO-70B-e2 50.81 100% 1.25%
IPL-IPO-70B-e3 52.13 100% 2.40%
Table 7: Analysis regards to details of IPL. The “-eN” notation refers to the “N”-th round of iteration in IPL training. The selection refers to the data proportion that be updated with newly generated prompts. The first epoch is for warm-up, therefore there are not scores.
Appendix D Entire fine-tuning results
We present entire fine-tuning results of our various experiments in Table 6. Apart from conclusions reported in Section 3.2.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected. However, all SFT models perform worse than the naive baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe that 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend, exhibiting a steeper upward growth of 2% per epoch. We suppose the regularized design in Equation 10 contributes to stable improvements.
6 BBH Tasks following PromptAgent Work Wang et al. (2023a)
Downstream LLM Optimizer Penguins Geometry Epistemic Object Counting Temporal Causal Judgment Average
Naive 0.436 0.536 0.611 0.318 0.362 0.547 0.468
APE 0.402 0.499 0.680 0.347 0.440 0.583 0.492
Llama2-7B PromptAgent 0.426 0.544 0.775 0.312 0.643 0.547 0.541
Touvron et al. (2023) GPT-4 0.444 0.578 0.640 0.392 0.536 0.598 0.531
FIPO 0.458 0.542 0.754 0.350 0.683 0.616 0.567
Naive 0.336 0.089 0.581 0.690 0.155 0.832 0.447
APE 0.313 0.400 0.627 0.525 0.189 0.882 0.489
Tulu2-7B PromptAgent 0.470 0.350 0.691 0.577 0.176 0.882 0.524
Ivison et al. (2023) GPT-4 0.542 0.397 0.603 0.674 0.176 0.888 0.547
FIPO 0.525 0.415 0.644 0.637 0.210 0.891 0.554
Naive 0.752 0.460 0.720 0.499 0.912 0.651 0.666
APE 0.740 0.533 0.713 0.530 0.897 0.643 0.676
Llama3-70B-Instruct PromptAgent 0.748 0.501 0.667 0.524 0.946 0.687 0.687
Dubey et al. (2024) GPT-4 0.756 0.478 0.726 0.472 0.912 0.644 0.665
FIPO 0.765 0.503 0.678 0.539 0.923 0.716 0.692
Naive 0.593 0.540 0.698 0.602 0.879 0.767 0.680
APE 0.614 0.496 0.718 0.617 0.888 0.790 0.687
Qwen2-72B-Instruct PromptAgent 0.658 0.533 0.717 0.622 0.898 0.779 0.722
Yang et al. (2024) GPT-4 0.630 0.540 0.700 0.642 0.876 0.759 0.691
FIPO 0.617 0.522 0.792 0.591 0.836 0.768 0.731
Naive 0.595 0.227 0.452 0.612 0.720 0.470 0.513
APE 0.747 0.490 0.708 0.716 0.856 0.570 0.681
GPT-3.5-turbo PromptAgent 0.797 0.670 0.806 0.860 0.934 0.670 0.790
(In-box Tester) GPT-4 0.632 0.589 0.840 0.647 0.821 0.550 0.680
FIPO 0.614 0.738 0.865 0.660 0.756 0.759 0.732
Naive 0.859 0.472 0.820 0.680 0.990 0.750 0.762
APE 0.862 0.691 0.844 0.650 0.988 0.749 0.797
GPT-4o PromptAgent 0.855 0.750 0.886 0.733 0.982 0.712 0.820
GPT-4 0.840 0.737 0.925 0.692 0.982 0.700 0.813
FIPO 0.876 0.810 0.891 0.684 0.965 0.840 0.844
Table 8: Comparison between FIPO optimizer, previous methods and GPT-4’s prompt optimization.
Appendix E Specific ablation discussion on BBH
We notice that the BBH dataset behaves significantly different from other datasets in the ablation experiments, as shown by the last lines of Table 6. The corrupted optimizer IPL-IPO-gen and IPL-IPO-partial gain suboptimal performances across all benchmarks, except BBH. We suppose this may be because the BBH contains a large number of symbolic reasoning tasks, and this unique pattern is not found in other testing datasets.
For instance, one of the typical subtask in BBH is “Name a Geometry”, which asks to “name geometric shapes from their SVG paths”. A typical example is listed here: “This SVG path element <path d="M 59.43,52.76 L 75.49,27.45 L 54.92,4.40 M 54.92,4.40 L 23.70,7.77 L 15.15,42.15 L 34.51,57.44 L 59.43,52.76"/> draws a ”. Therefore, we suspect that when further using discrete natural language instructions to diversify our training data, this may affect subtasks in BBH that involve specific symbolic reasoning patterns.
Nevertheless, when examining the six representative subtasks in BBH, FIPO still achieves a clear advantage over other leading prompt optimization methods, as shown by the Table 8.
Appendix F Detailed results of 6 BBH subsets
We provide the performance of various optimizers on different downstream testers, including the evaluation results of 6 BBH tasks in Table 8. Here are some key points of the result analysis:
  • Overall Performance. The FIPO optimizer performs well in most combinations, usually leading in each task and average score. PromptAgent and GPT-4 optimizers perform well on some specific tasks, but are slightly inferior to FIPO in overall average scores.
  • Optimizer Comparison. When comparing different optimizers, FIPO outperforms other methods across multiple downstream LLMs. While PromptAgent occasionally achieves high scores in individual tasks, its average performance is generally lower than FIPO. Similarly, the GPT-4 optimizer shows competitive results in certain scenarios but does not maintain the same level of consistency as FIPO. On the other hand, the Naive prompt and APE optimizer often lag behind in most tasks. For example, on GPT-4o, APE’s average score is 0.797, lower than FIPO’s 0.844.
  • Task-specific Performance. On a task level, FIPO excels in many tasks. For instance, in the "Temporal" and "Causal Judgment" tasks under Llama2-7B, it achieved the highest scores of 0.683 and 0.616, respectively. It also secured the top scores of 0.525 and 0.891 in the "Penguins" and "Causal Judgment" tasks under Tulu2-7B. For GPT-4o, FIPO achieved the highest scores in the "Geometry" and "Causal Judgment" tasks, with 0.810 and 0.840, respectively. Similarly, in the "Epistemic" task under Qwen2-72B-Instruct, FIPO led with a score of 0.792. Overall, FIPO optimizer demonstrates balanced and strong performance across various tasks.
In summary, the FIPO optimizer performed best on a variety of downstream LLMs, especially on GPT-4o and Llama3-70B-Instruct, where its stability and efficiency were remarkable. The PromptAgent and GPT-4 optimizers also performed well on specific tasks, but were not as comprehensive as FIPO overall. The Naive prompt and APE optimizer performed relatively poorly.
Appendix G Cost comparison: FIPO vs. others
In terms of overall cost against other aforementioned optimization approaches, shown in Table 9, building FIPO optimizer will have \$300 one-time collecting cost and 10 hours training cost, which equals to 75 usages of GPT-4. However, once FIPO models are prepared, we save a lot of time as the optimization can be done locally. FIPO leads to less total costs when optimizing over massive prompts.
In specific, we explain the cost number in Table 9 column by column. The first column refers to the cost of constructing data. Since APE, PromptAgent and directly using GPT-4 are all training-free methods, they do not have the cost of constructing training data. Similarly, the second column refers to the time cost of training, and only our local fine-tuned FIPO optimizer covers that. As for the inference cost in the third and fourth columns, we report the inference cost of the entire representative “penguins_in_a_table” subtask in BBH, which includes 149 test samples. In particular, it worth to note that the inference cost of PromptAgent is consistent with its official report [^6] https://github.com/XinyuanWangCS/PromptAgent/. In addition, as the only local optimizer, FIPO also includes a one-time inference deployment cost. However, in our practice, since the time for renting a server and deploying a model does not exceed 3 minutes in total, this part of the cost is basically negligible.
Appendix H Overwhelmed Cheating Notes
We provide a proportion visualization of analyzing the possible over-optimized prompts in each testing benchmark, manually checking with 256 random samples, in Figure 6. The results show that mathematical questions in GSM8K and BBH are more likely to be added cheating notes of final answers, while other types of testing data receive moderate prompt optimization. Nevertheless, even taking into account this factor, the conclusion that our local optimizer improves the overall performance of general downstream generators by optimizing prompts reasonably does not change (§ 3.2.1).
Inference Per Test
Dataset Training Fee Time
APE \$0 \$0 \$5 2h
PromptAgent \$0 \$0 \$5 2h
GPT-4 \$0 \$0 \$4 1h
FIPO (IPL-IPO-70B) \$300 \$60 \$0 30s
Table 9: Cost report of different methods.
Figure 6: Analysis of overwhelmed cheating notes on 256 random optimized prompts for each benchmark.
Appendix I Employed Meta Prompts
We report all employed meta prompts in this section as reference for potential future researches. Please refer to Table 10,  11 and  12 for data collection meta prompt, data diversification meta prompt, and discrimination meta prompt, respectively.
Data collection meta prompt
You are an expert of prompt optimization.
Sliver Prompt:
SP
Sliver Response:
SR
Golden Response:
GR
Task Introduction:
Based on the Silver Prompt, optional Silver Response and optional Golden Response, perform the following actions:
1 – The optional Sliver Response was generated by an AI based on the Silver Prompt. Please help modify the Silver Prompt to Golden Prompt that can obtain a more correct response, in reference to the optional Golden Response.
2 - When building the Golden Prompt, you can consider several aspects, such as: (1) A roleplay leading sentence to adapt the AI to the task-specific scenario; (2) Details of task characteristics, for instance, the task could be a question answering task, a dialogue task, or a summarization task, etc; (3) Further clarification of the task information, especially some ambiguous terms; (4) A more detailed solution guidance, such as step-by-step plans, handlings of exceptions, special priorities or constraints, etc; (5) Any specific requirements for the response, such as the length, the format, the style, the tone, the language, etc.
3 - Show me only the Golden Prompt, do not contain any other content.
Golden Prompt:
Table 10: The meta prompt used to harness data from GPT-3.5-turbo and GPT-4 APIs. During harnessing, we replace placeholders “SP”, “SR” and “GR” with actual naive prompt $\bm{x^{n}}$, naive response $\bm{\hat{y}^{n}}$ and ground truth response $\bm{y}^{n}$ from each seed data sample, respectively.
Data diversification meta prompt
You are an expert of prompt optimization.
Sliver Prompt:
SP
<Optional Responses>
The optional Sliver Response was generated by an AI based on the Silver Prompt. Please help modify the Silver Prompt to Golden Prompt that can obtain a more correct response, in reference to the optional Golden Response. The Golden Prompt should be instructive, concise and strictly faithful to any factual information in the Silver Prompt. The length of the Golden Prompt should be less than GN words. Only give me the content of Golden Prompt, do not contain any other information (e.g., your response of the Golden Prompt, any postfix like “Golden Prompt”, etc.).
Flexible meta prompt for optional naive response
Sliver Response:
SR
Flexible meta prompt for optional ground truth response
Golden Response:
GR
Table 11: The meta prompt used for data diversification. During diversification, we replace placeholders “SP”, “SR” and “GR” with actual naive prompt $\bm{x^{n}}$, naive response $\bm{\hat{y}^{n}}$ and ground truth response $\bm{y}^{n}$, respectively. Meanwhile, flexible prompts of optional naive response and ground truth response are inserted by probability as introduced in Figure 3 and  5. Moreover, we add length suggestion with “GN” signal, using the length of chosenl optimized prompt label $\bm{x^{o+}}$.
Discrimination meta prompt
You are an expert of prompt discrimination.
Raw Prompt:
RP
New Prompt A:
PA
New Prompt B:
PB
Golden Response:
GR
New Prompt A and New Prompt B are optimized from Raw Prompt. Please judge which prompt is more loyal to the factual information of Raw Prompt, and is more desirable for an AI to generate the Golden Response. Only answer with A or B.
Table 12: The meta prompt used for creating discrimination data and application of discrimination during IPL training. The creation of discrimination data has been introduced in the right part of Figure 5. In terms of inner discrimination of IPL, we replace placeholders “RP”, “PA”, “PB” and “GR” with naive prompt $\bm{x^{n}}$, naive prompt $\bm{x^{n}}$, newly optimized prompt $\bm{x^{n+}}$ and ground truth response $\bm{y}^{n}$, respectively.